For the chemical equation

 SO2(g) + NO2(g) ⇌ SO3(g) + NO(g)

 the equilibrium constant at a certain temperature is 2.80. At this temperature, calculate the number of moles of NO2(g) that must be added to 3.36 mol SO2(g) in order to form 1.40 mol SO3(g) at equilibrium.
Answer #1

Let mol of NO2 added be n

SO2(g) + NO2(g) <-> SO3(g) + NO(g)

3.36 n 0 0 (initial)

3.36-x n-x x x (at equilibrium)

Mol of SO3 formed = x = 1.40 mol

Kc = [SO3][NO]/[SO2][NO2]

2.80 = x*x / (3.36-x)(n-x)

2.80 = (1.40*1.40) / (3.36-1.40)(n-1.40)

2.80 = 1/(n-1.40)

2.80*n - 3.92 = 1

2.80*n = 4.92

n = 1.76 mol

Answer: 1.76 mol